aloevera:

What is the work done by a 20 newton force applied at an angle of 45.0° to move a box a horizontal distance of 40 meters? A. 8.0 × 102 joules B. 9.0 × 102 joules C. 5.6 × 102 joules D. 3.6 × 102 joules

dude:

To calculate the work needed use the equation \(W=F\times D\times cos(\theta)\) F= Force D = Distance \(\theta\)= Degree/angle We know that the for is 20 N, distance is 40m, and the angle at which it was dropped was 45º

dude:

Oh also, once you calculate it, convert the answer to scientific notation